# Enhancing Real-World Complex Network Representations with Hyperedge Augmentation

### v1

@@ Appendix F Ethics Statement
GCN+HyperAug(GAT) 59.0 $\pm$ 0.2 58.3 $\pm$ 0.2 58.1 $\pm$ 0.4 58.4 $\pm$ 0.2
GCN+HyperAug(GraphSAGE) 60.1 $\pm$ 0.2 64.1 $\pm$ 0.2 62.1 $\pm$ 0.2 62.1 $\pm$ 0.2
GAT+HyperAug(GraphSAGE) 61.8 $\pm$ 0.5 63.2 $\pm$ 0.7 62.4 $\pm$ 1.9 61.2 $\pm$ 1.1
HyperConv+HyperAug(ED-HNN) 58.9 $\pm$ 1.4 61.5 $\pm$ 1.0 62.3 $\pm$ 2.1 61.1 $\pm$ 0.3
GCN+HyperAug(HyperConv) 60.4 $\pm$ 0.4 61.4 $\pm$ 0.6 60.5 $\pm$ 0.3 60.9 $\pm$ 0.4
GCN+HyperAug(ED-HNN) 61.5 $\pm$ 1.0 63.7 $\pm$ 0.9 63.4 $\pm$ 2.0 62.4 $\pm$ 1.3
GAT+HyperAug(HyperConv) 62.5 $\pm$ 1.2 61.8 $\pm$ 0.3 64.2 $\pm$ 1.6 64.0 $\pm$ 0.3
GAT+HyperAug(ED-HNN) 62.3 $\pm$ 1.2 61.5 $\pm$ 1.2 64.5 $\pm$ 1.9 63.8 $\pm$ 1.8
GraphSAGE+HyperAug(HyperConv) 62.9 $\pm$ 0.5 66.4 $\pm$ 0.2 66.3 $\pm$ 1.5 64.8 $\pm$ 0.4
GraphSAGE+HyperAug(ED-HNN) 62.6 $\pm$ 0.3 64.5 $\pm$ 0.1 64.7 $\pm$ 0.7 63.9 $\pm$ 0.2
Table 13. Accuracy (%) of the HyperAug variants and vanilla GNN baselines on the citation datasets. *The results for GCN and GAT are directly taken from the GAT paper (Veličković et al., 2018). **The results for HyperConv and ED-HNN are directly taken from the ED-HNN paper (Wang et al., 2023).
@@
GCN+HyperAug(GAT) 86.9 $\pm$ 0.6 86.9 $\pm$ 0.6 87.0 $\pm$ 1.0
GCN+HyperAug(GraphSAGE) 87.4 $\pm$ 1.4 87.4 $\pm$ 1.4 87.4 $\pm$ 0.6
GAT+HyperAug(GraphSAGE) 85.8 $\pm$ 0.0 85.8 $\pm$ 0.0 87.3 $\pm$ 1.0
HyperConv+HyperAug(ED-HNN) 83.0 $\pm$ 1.0 85.2 $\pm$ 1.4 87.4 $\pm$ 0.1
GCN+HyperAug(HyperConv) 86.9 $\pm$ 1.7 87.5 $\pm$ 0.9 87.3 $\pm$ 0.4
GCN+HyperAug(ED-HNN) 86.6 $\pm$ 1.4 87.0 $\pm$ 0.8 89.0 $\pm$ 0.2
GAT+HyperAug(HyperConv) 86.7 $\pm$ 0.4 86.0 $\pm$ 2.2 86.6 $\pm$ 0.7
GAT+HyperAug(ED-HNN) 87.0 $\pm$ 1.3 86.7 $\pm$ 0.0 89.0 $\pm$ 0.6
GraphSAGE+HyperAug(HyperConv) 87.2 $\pm$ 1.8 86.3 $\pm$ 0.8 87.3 $\pm$ 0.4
GraphSAGE+HyperAug(ED-HNN) 87.4 $\pm$ 1.0 87.7 $\pm$ 0.8 88.8 $\pm$ 1.1
Table 14. Accuracy (%) of the HyperAug variants and vanilla GNN baselines on the Amazon datasets.
@@
GCN+HyperAug(GAT) 91.3 $\pm$ 0.1 71.1 $\pm$ 0.5
GCN+HyperAug(GraphSAGE) 93.0 $\pm$ 0.0 78.6 $\pm$ 0.6
GAT+HyperAug(GraphSAGE) 96.8 $\pm$ 0.6 78.0 $\pm$ 0.1
HyperConv+HyperAug(ED-HNN) 97.8 $\pm$ 0.1 80.7 $\pm$ 0.6
GCN+HyperAug(HyperConv) 96.8 $\pm$ 0.5 69.8 $\pm$ 0.5
GCN+HyperAug(ED-HNN) 98.1 $\pm$ 0.7 80.5 $\pm$ 0.6
GAT+HyperAug(HyperConv) 96.8 $\pm$ 0.6 71.5 $\pm$ 0.3
GAT+HyperAug(ED-HNN) 98.2 $\pm$ 0.1 80.9 $\pm$ 0.3
GraphSAGE+HyperAug(HyperConv) 98.2 $\pm$ 0.7 79.3 $\pm$ 0.8
GraphSAGE+HyperAug(ED-HNN) 98.1 $\pm$ 0.1 80.9 $\pm$ 0.7
Figure 4. Node prediction performances of HyperConv combined with different GraphSAINT sampling techniques on 21 virtual hypergraphs produced by HyperAug, excluding MUSAE-GitHub and MUSAE-Facebook.

### v2

@@ Appendix E Ethics Statement
GCN+TopoAug(GAT) 59.0 $\pm$ 0.2 58.3 $\pm$ 0.2 58.1 $\pm$ 0.4 58.4 $\pm$ 0.2
GCN+TopoAug(GraphSAGE) 60.1 $\pm$ 0.2 64.1 $\pm$ 0.2 62.1 $\pm$ 0.2 62.1 $\pm$ 0.2
GAT+TopoAug(GraphSAGE) 61.8 $\pm$ 0.5 63.2 $\pm$ 0.7 62.4 $\pm$ 1.9 61.2 $\pm$ 1.1
HyperConv+TopoAug(ED-HNN) 58.9 $\pm$ 1.4 61.5 $\pm$ 1.0 62.3 $\pm$ 2.1 61.1 $\pm$ 0.3
GCN+TopoAug(HyperConv) 60.4 $\pm$ 0.4 61.4 $\pm$ 0.6 60.5 $\pm$ 0.3 60.9 $\pm$ 0.4
GCN+TopoAug(ED-HNN) 61.5 $\pm$ 1.0 63.7 $\pm$ 0.9 63.4 $\pm$ 2.0 62.4 $\pm$ 1.3
GAT+TopoAug(HyperConv) 62.5 $\pm$ 1.2 61.8 $\pm$ 0.3 64.2 $\pm$ 1.6 64.0 $\pm$ 0.3
GAT+TopoAug(ED-HNN) 62.3 $\pm$ 1.2 61.5 $\pm$ 1.2 64.5 $\pm$ 1.9 63.8 $\pm$ 1.8
GraphSAGE+TopoAug(HyperConv) 62.9 $\pm$ 0.5 66.4 $\pm$ 0.2 66.3 $\pm$ 1.5 64.8 $\pm$ 0.4
GraphSAGE+TopoAug(ED-HNN) 62.6 $\pm$ 0.3 64.5 $\pm$ 0.1 64.7 $\pm$ 0.7 63.9 $\pm$ 0.2
Table 12: Accuracy (%) of the TopoAug variants and vanilla GNN baselines on the citation datasets. *The results for GCN and GAT are directly taken from the GAT paper [33]. **The results for HyperConv and ED-HNN are directly taken from the ED-HNN paper [34].
@@
GCN+TopoAug(GAT) 86.9 $\pm$ 0.6 86.9 $\pm$ 0.6 87.0 $\pm$ 1.0
GCN+TopoAug(GraphSAGE) 87.4 $\pm$ 1.4 87.4 $\pm$ 1.4 87.4 $\pm$ 0.6
GAT+TopoAug(GraphSAGE) 85.8 $\pm$ 0.0 85.8 $\pm$ 0.0 87.3 $\pm$ 1.0
HyperConv+TopoAug(ED-HNN) 83.0 $\pm$ 1.0 85.2 $\pm$ 1.4 87.4 $\pm$ 0.1
GCN+TopoAug(HyperConv) 86.9 $\pm$ 1.7 87.5 $\pm$ 0.9 87.3 $\pm$ 0.4
GCN+TopoAug(ED-HNN) 86.6 $\pm$ 1.4 87.0 $\pm$ 0.8 89.0 $\pm$ 0.2
GAT+TopoAug(HyperConv) 86.7 $\pm$ 0.4 86.0 $\pm$ 2.2 86.6 $\pm$ 0.7
GAT+TopoAug(ED-HNN) 87.0 $\pm$ 1.3 86.7 $\pm$ 0.0 89.0 $\pm$ 0.6
GraphSAGE+TopoAug(HyperConv) 87.2 $\pm$ 1.8 86.3 $\pm$ 0.8 87.3 $\pm$ 0.4
GraphSAGE+TopoAug(ED-HNN) 87.4 $\pm$ 1.0 87.7 $\pm$ 0.8 88.8 $\pm$ 1.1
Table 13: Accuracy (%) of the TopoAug variants and vanilla GNN baselines on the Amazon datasets.
@@
GCN+TopoAug(GAT) 91.3 $\pm$ 0.1 71.1 $\pm$ 0.5
GCN+TopoAug(GraphSAGE) 93.0 $\pm$ 0.0 78.6 $\pm$ 0.6
GAT+TopoAug(GraphSAGE) 96.8 $\pm$ 0.6 78.0 $\pm$ 0.1
HyperConv+TopoAug(ED-HNN) 97.8 $\pm$ 0.1 80.7 $\pm$ 0.6
GCN+TopoAug(HyperConv) 96.8 $\pm$ 0.5 69.8 $\pm$ 0.5
GCN+TopoAug(ED-HNN) 98.1 $\pm$ 0.7 80.5 $\pm$ 0.6
GAT+TopoAug(HyperConv) 96.8 $\pm$ 0.6 71.5 $\pm$ 0.3
GAT+TopoAug(ED-HNN) 98.2 $\pm$ 0.1 80.9 $\pm$ 0.3
GraphSAGE+TopoAug(HyperConv) 98.2 $\pm$ 0.7 79.3 $\pm$ 0.8
GraphSAGE+TopoAug(ED-HNN) 98.1 $\pm$ 0.1 80.9 $\pm$ 0.7